rolinsedu/webopps/assessment/tabe lauech jnplcouse assessment id 19149 1Bcourse d22 701 1Bcontent is 533152 1astepsnul Consider the following production line consisting of 5 sectors and 12 machines. The individual machine capacity and the number of machines in each sector are given below Sector 1 Sector 2 Sector 3 Sector 4 Sector 5 21 3 4 5 2 2 3 Individual machines2 hrsjunit 3 hrs/unit 7 hrsunit 5 hrs/unit 1.5 hrsJunit capacity What is the capacity of the line? a 71 units per hour b. 66 units per hour C.36 units per hour d. 50 units per hour
Answer #1

Answer:

  • Option d: 0.50 units per hour

Explanation:

  • Capacity of each sector = no. of machine / capacity given in question
  • Following are the capacity of each sector :
    • ​1 : 1/2 = 0.5 unit / hr
    • 2 : 2/3 = 0.66 unit /hr
    • 3 : 5 /7 = 0.71 unit / hr
    • 4 : 3 /5 = 0.6 unit /hr
    • 5 : 1/ 1.5 = 0.66 units / hr
  • Hence the minimum capacity is of sector 1 , so the process capacity is the capacity of bottle neck = 0.5 unit /hr